OpenStudy (theloshua):

4z^3 - 8z^2 = 0 solve for x

Vocaloid (vocaloid):

I think you mean "solve for z"? start by factoring

OpenStudy (anonymous):

can't wait to see the answer to this

OpenStudy (theloshua):

Z

OpenStudy (theloshua):

I meant z, sorry :p

Vocaloid (vocaloid):

anyway, start by identifying the greatest common factor of 4z^2 and 8z^2

Vocaloid (vocaloid):

*4z^3 and 8z^2

OpenStudy (theloshua):

4z^2(z - 2) = 0 right?

Vocaloid (vocaloid):

excellent! now we split it into two parts 4z^2 = 0 and (z-2) = 0 solve for z in both instances

OpenStudy (theloshua):

z = 2 z = 0 :D

OpenStudy (theloshua):

thnx